A chemist started with 26.60 mL of an aqueous glucose solution and added 97.10 mL of water followed by stirring. The resultant glucose concentration was 0.47 mol L-1. What was the concentration of glucose in the first solution (in mol L-1)?
1
Expert's answer
2019-01-14T04:27:11-0500

Answer:

V1=26.6 ml = 0.0266 L

V2=97.1 ml =0.0971 L

Vtotal=0.0266+0.0971 = 0.1237 L

CM=0.47 M

CMfirst=?

Formula:

CM=n/V; n=CM x V;

n=0.47 x 0.1237 = 0.058139 mol.

CMfirst=0.058139 / 0.0266 =2.1856 M